# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ Abstract
Standard multiple testing procedures are designed to report a list of discoveries, or suspected false null hypotheses, given the hypotheses’ p-values or test scores. Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score. Specifically, such so-called “side information” can be leveraged to improve the separation between true and false nulls along additional “dimensions” thereby increasing the overall sensitivity. In line with this idea, we develop RESET (REScoring via Estimating and Training) which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side-information while maintaining finite-sample error rate control. Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible to a range of multiple testing scenarios without the need for the user to select the appropriate one. We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) is able to uniquely achieve finite sample FDR or FDP control, depending on the user’s preference.
* Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
@@ 1 Introduction
Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$. As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment. In this case, the $i$th null hypothesis $H_{i}$ states that “the $i$th gene exhibits no change in gene expression level”. This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis. The collection of p-values then undergoes a filtering process that rejects a subset of these null hypotheses that have sufficiently small p-values subject to a type-1 error rate control to minimize the number of misreported true null hypotheses.
@@
The most common choice of error rate in this setup is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries [2]. In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold. Alternatively, we might be interested in controlling the FDP rather than its expectation in the following sense. Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$. Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$ which is not guaranteed by FDR control [39, 27, 19, 37, 16].
@@
Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution, and (b) it can be computationally intensive. A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values. Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution. The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
@@
Those p-values are too crude to be used in typical p-value based filters. Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$. The function $f$ needs to be symmetric in its arguments and should be large if $Z_{i}$ is large, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or if $Z_{i}$ is larger than its counterpart $\tilde{Z}_{i}$, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$ [^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$.. For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric. Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses. Using this fact, the scores and labels $\{(W_{i},L_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
@@
More recently, research has been focused on improving these filtering methods by leveraging side-information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses. These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side-information is used to weight or rescore the hypotheses according to how confident we believe each hypothesis is a false null. Indeed, there are several such tools that augment p-values with side-information [25, 34, 33, 36, 50, 35]. When it comes to using side-information in the competition framework, we are aware of only one method with proper error-rate control along with a precise implementation, namely the recently published Adaptive Knockoffs [43] (AdaPT [33] offers a meta-procedure, but without a precise implementation).
@@
In this paper, we offer a novel approach for multiple testing with side-information. The idea is to randomly split a suspected subsets of the true nulls into two sets: one set is used for training a semi-supervised learning model to discriminate the true and false null hypotheses, while the second is reserved for estimating the number of false discoveries. Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training. Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses. Moreover, it works the same in the competition and p-value context and allows the user to choose between the choice of FDR or FDP control. We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
@@ 2 Background: competition-based multiple testing
Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e. long chains of amino acids, are present in a biological sample [17]. Direct protein identification is challenging and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead. Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum. Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
@@
This is done by searching each spectrum against a database of candidate peptides called targets to find its optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match. Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect. Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that “the $i$th target does not belong to the sample”. In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence. Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
@@
As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated to the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample. Similarly we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide. For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else. Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [41].
@@
The statistics community independently formalized this idea in the context of variable selection where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$. The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff. In this paper, we consider the model-X version of the knockoff filter [7] which makes the following assumptions [^2] However, our results are equally applicable to the fixed-X model as well [1].. First, the joint distribution of the variables, $X$, is assumed to be known. Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d from the joint distribution $F_{XY}$. Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly based on $X$, requiring that they satisfy the following conditions:
@@
Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}:=W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$. The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much. Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$, however as it will become clear in our later presentation of RESET, it is instructive to separate them. In this paper, we consider the Lasso Coefficent Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [48]. Specifically, cross-validation is used to to determine the $\lambda$ penalty when regresson on the augmented design matrix $[X,\tilde{X}]$. We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs. The scores and labels are then defined as:
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
For the knockoff filter, Assumption 1 follows from the knockoff construction while in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and morever, this happens independently of all other pairs of scores. This assumption is commonly applied in the literature, with several analyses validating this claim empirically [38, 21].
Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e. decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$. Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways. The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1] which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\}.$ (2)
SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$. In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short. As such, we will refer to the method as SSS+ in the variable selection setting, and TDC in the peptide detection setting.
Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[39]. FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order. Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$ which is determined by $\alpha$ and $\gamma$. Each comparison is made in a ‘stepdown’ fashion, i.e, if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves onto the next comparison at $i+1$. For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$. Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP.

### v2

# A semi-supervised framework for diverse multiple hypothesis testing scenarios

@@ Abstract
Standard multiple testing procedures are designed to report a list of discoveries, or suspected false null hypotheses, given the hypotheses’ p-values or test scores. Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score. Specifically, such so-called side information can be leveraged to improve the separation between true and false nulls along additional dimensions thereby increasing the overall sensitivity. In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite-sample error rate control. Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible with a range of multiple testing scenarios without the need for the user to select the appropriate one. We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery proportion control, depending on the user’s preference.
Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
@@ 1 Introduction
Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$. As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment. In this case, the $i$th null hypothesis $H_{i}$ states that ‘the $i$th gene exhibits no change in gene expression level’. This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis. The collection of p-values then undergoes a filtering process that rejects a subset of these null hypotheses that have sufficiently small p-values subject to a type-1 error rate control to minimize the number of misreported true null hypotheses.
@@
The most common choice of error rate in this setup is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries [2]. In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold. Alternatively, we might be interested in controlling the FDP rather than its expectation in the following sense. Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$. Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control [40, 27, 19, 38, 16, 20].
@@
Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution and (b) it can be computationally intensive. A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values [1, 21]. Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution. The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
@@
These p-values are too crude to be used in typical p-value based filters. Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$. The function $f$ needs to be symmetric in its arguments and should increase with the values of $Z_{i}$ and $\tilde{Z}_{i}$, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or with the discrepancy between the two, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$. [^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$. For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric. Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses. Using this fact, the labels and scores $\{(L_{i},W_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
@@
More recently, research has been focused on improving these filtering methods by leveraging side information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses. These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side information is used to weight or rescore the hypotheses according to how confidently we believe that each hypothesis is a false null. Indeed, there are several such tools that augment p-values with side information [25, 34, 33, 37, 51, 35]. When it comes to using side information in the competition framework, we are aware of only one fully specified method with rigorous error-rate control, which is the recently published Adaptive Knockoffs [44] (AdaPT [33] only offers a meta-procedure).
@@
In this paper, we offer a novel approach for multiple testing with side information. The idea is to randomly split a set of suspected true nulls into two sets, where one set is used for training a semi-supervised learning model to discriminate between the true and false null hypotheses, and the other for estimating the number of false discoveries. Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training. Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses. Moreover, it is equally applicable to the competition and p-value context and allows the user to choose between FDR or FDP control. We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
@@ 2 Background: competition-based multiple testing
Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample [17]. Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead. Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum. Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
@@
This is done by searching each spectrum against a database of candidate peptides called targets to find the spectrum’s optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match. Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect. Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that ‘the $i$th target is not present in the sample’. In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence. Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
@@
As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample. Similarly, we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide. For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else. Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [42].
@@
Subsequently, the statistics community independently formalized this idea in the context of variable selection, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$. The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff. In this paper, we consider the model-X version of the knockoff filter [7] which makes the following assumptions [^2] However, our results are equally applicable to the fixed-X model as well [1].. First, the joint distribution of the variables, $X$, is assumed to be known. Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d. from the joint distribution $F_{XY}$. Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly subject to the following conditions:
@@
Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}:=W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$. The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much. Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, as will become clear in our later presentation of RESET, it is instructive to separate them.
In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [49]. Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $[X,\tilde{X}]$. We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs. The scores and labels are then defined as
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
For the knockoff filter, Assumption 1 follows from the knockoff construction, whereas in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and moreover, this happens independently of all other pairs of scores. This assumption is commonly applied in the literature, with several analyses validating this claim empirically [39, 21].
Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e., decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$. Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways. The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1], which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\left\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\right\}.$ (2)
SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$. In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short. As such, we will refer to the method as TDC in the peptide detection setting, and SSS+ in the variable selection setting and more generally.
Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[40]. FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order. Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$. Each comparison is made in a ‘stepdown’ fashion, i.e., if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves on to the next comparison at $i+1$. For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$. Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP: